How do I find the value of k so that the line passing through has the given slope m in the problem
(-k+1,3), (3,k+3), m=3

Answer by lwsshak3(11628)

How do I find the value of k so that the line passing through has the given slope m in the problem
(-k+1,3), (3,k+3), m=3
**
slope=∆y/∆x=(y1-y2)/(x1-x2)
3=(k+3-3)/(3+k-1)=k/k+2
k=3k+6
-2k=6
k=-3